Perfect Pet Collar Company makes custom leather pet collars. The company expects each collar to require 1.80 feet of leather and predicts leather will cost $3.10 per foot. Suppose Perfect Pet made 75 collars during February. For these 75 collars, the company actually averaged 2.10 feet of leather per collar and paid $2.60 per foot.

1. Calculate the standard direct materials cost per unit.

2. Without performing any calculations, determine whether the direct materials price variance will be favorable or unfavorable.

3. Without performing any calculations, determine whether the direct materials quantity variance will be favorable or unfavorable.

6. Calculate the direct materials price and quantity variances.
Answer #1

a) Standard direct material cost per unit = 1.80*3.1 = 5.58

2) Direct material price variance will be favorable because actual cost is less the standard cost

3) Direct material quantity variance will be unfavorable because actual quantity would be higher than standard qty

4) Material price variance = (3.10-2.60)*157.50 = 78.75 F

Material quantity variance = (1.80-2.10*75)*3.10 = 69.75 U